# Efficient Long Sequential Low-rank Adaptive Attention for Click-through rate Prediction

### v2

@@ 3.1. Long Sequence Compression
In this equation, $E_{s}^{T}$ is the transposed matrix of $E_{s}$, and $AttScore(E_{s},$ $E_{t})\in\mathbb{R}^{L\times 1}$ represents the attention score designed to the behavior sequence of length $L$, which is a critical step in CTR prediction. However, the calculation of target attention becomes a bottleneck during inference due to the constraints posed by the lengths of long-term sequences.
@@
Here, $W_{Comp}\in\mathbb{R}^{L\times r}$ and $W_{Decopm}\in\mathbb{R}^{r\times L}$, where $L$ is the original sequence length and $r$ is the reduced rank. $E_{comp}=E_{s}^{T}W_{Comp}\in\mathbb{R}^{d\times r}$ can be seen as a compressed representation of the behavior sequence. This operation helps to preserve essential information while handling the compressive aspect of the sequence. We can express target attention computation like implementation of DIN target attention(Zhou et al., 2018) as:
@@
(9) $\begin{split}LREA(E_{s}^{T},E_{t})=E_{S}^{T}W_{Decomp}^{T}\phi\bigg{(}[E_{t}\oplus W_{comp}^{T}\oplus(E_{t}\odot W_{comp}^{T})]W_{in}\bigg{)}W_{o}\end{split}$
The $W_{Decomp}$ could be absorbed by $E_{auxabsorb}=E_{S}^{T}W_{Decomp}^{T}\in\in\mathbb{R}^{d\times r}$, just like $E_{comp}$, It compression the L-dimension of $E_{s}$ to a low-rank $r$ Finally the output of LREA will be concatenated with all other features’ embedding and then fed into MLP to predict $\hat{y}$.

### v3

@@ 3.1. Long Sequence Compression
In this equation, $E_{s}^{T}$ is the transposed matrix of $E_{s}$, and $AttScore(E_{s},$ $E_{t})\in\mathbb{R}^{L\times 1}$ represents the attention score correspond to each item in the behavior sequence of length $L$, which is a critical step in CTR prediction. However, the calculation of target attention becomes a bottleneck during inference due to the constraints posed by the lengths of long-term sequences.
@@
Here, $W_{Comp}\in\mathbb{R}^{L\times r}$ and $W_{Decopm}\in\mathbb{R}^{r\times L}$, where $L$ is the original sequence length and $r$ is the reduced rank. $E_{comp}=E_{s}^{T}W_{Comp}\in\mathbb{R}^{d\times r}$ can be seen as a compressed representation of the behavior sequence. This operation helps to preserve essential information while handling the compressive aspect of the sequence. We can express attention score computation like implementation of DIN target attention(Zhou et al., 2018) score as:
@@
(9) $\begin{split}LREA(E_{s}^{T},E_{t})=E_{S}^{T}W_{Decomp}^{T}\phi\bigg{(}[E_{t}\oplus E_{comp}^{T}\oplus(E_{t}\odot E_{comp}^{T})]W_{in}\bigg{)}W_{o}\end{split}$
The $W_{Decomp}$ could be absorbed by $E_{auxabsorb}=\bigg{(}E_{S}^{T}W_{Decomp}^{T}\bigg{)}_{(d\times r)}$, just like $E_{comp}$, It compression the L-dimension of $E_{s}$ to a low-rank $r$ Finally the output of LREA will be concatenated with all other features’ embedding and then fed into MLP to predict $\hat{y}$.